# A Diagnostic to Find and Help Combat Positivity Issues – with a Focus on Continuous Treatments

### v1

@@ B.1 Polynomial
B.1.1 Causal Dose Response Curve and Naive Shift Response Curve
The results for the CDRC and the naive shift response curve for the polynomial data are given in Figures 8(a) and 8(b).
@@ B.1.1 Causal Dose Response Curve and Naive Shift Response Curve
B.1.2 Threshold Response Curve and Conditional Shift Response Curve
The results for the threshold response curve and the conditional shift response curve for the polynomial data are given in Figures 9(a) and 9(b).
@@ B.1.2 Threshold Response Curve and Conditional Shift Response Curve
B.2 Piece-wise Linear
@@
B.2.1 Causal Dose Response Curve and Naive Shift Response Curve
The results for the CDRC and the naive shift response curve for the piece-wise linear data are given in Figures 10(a) and 10(b).
@@ B.2.1 Causal Dose Response Curve and Naive Shift Response Curve
B.2.2 Threshold Response Curve and Conditional Shift Response Curve
The results for the threshold response curve and the conditional shift response curve for the polynomial data are given in Figures 11(a) and 11(b).

### v3

@@ C.1 Polynomial
C.1.1 Causal Dose Response Curve and Naive Shift Response Curve
The results for the CDRC and the naive shift response curve for the polynomial data are given in Figures 9(a) and 9(b).
@@ C.1.1 Causal Dose Response Curve and Naive Shift Response Curve
C.1.2 Threshold Response Curve and Conditional Shift Response Curve
The results for the threshold response curve and the conditional shift response curve for the polynomial data are given in Figures 10(a) and 10(b).
@@ C.1.2 Threshold Response Curve and Conditional Shift Response Curve
C.2 Piece-wise Linear
@@
C.2.1 Causal Dose Response Curve and Naive Shift Response Curve
The results for the CDRC and the naive shift response curve for the piece-wise linear data are given in Figures 11(a) and 11(b).
@@ C.2.1 Causal Dose Response Curve and Naive Shift Response Curve
C.2.2 Threshold Response Curve and Conditional Shift Response Curve
The results for the threshold response curve and the conditional shift response curve for the polynomial data are given in Figures 12(a) and 12(b).